the kinetic energy of 3 moles of a gas is  (a)3/2RT  (b)3RT  (C)1/2RT  (D)2RT

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
E = 3/2 KT
 
Now
K= R/n
E= 3/2 * R/3 T
 Hence
E= 1/2 RT
 
Hence option C is correct
 
Regards
Arun (askIITians forum expert)